# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

### v1

@@ Appendix A Summary Tables of Scientific LLMs
CATTS Cachola et al. (2020) L 406M BART SciTLDR sequence to sequence paper summarization
@@
SciMult Zhang et al. (2023g) L+G 138M MoE MAG, classification, classification, link prediction,
@@
MAmmoTH Yue et al. (2023) L 7B, 13B, LLaMA-2 MathInstruct instruction tuning QA, MWP
@@
MetaMath Yu et al. (2023b) L 7B, 13B, LLaMA-2 MetaMathQA instruction tuning QA, MWP
@@
MathCoder Wang et al. (2023e) L 7B, 13B, LLaMA-2 MathCodeInstruct instruction tuning QA, MWP
@@
OVM Yu et al. (2023a) L 7B LLaMA-2 GSM8K supervised fine-tuning QA, MWP, quantitative reasoning
@@
MAmmoTH2 Yue et al. (2024) L 8B LLaMA-3 WebInstruct instruction tuning QA, MWP, quantitative reasoning
@@
     Liang et al. (2023b) VQ-VAE TabMWP sequence to sequence geometry problem solving
@@
Table-GPT Li et al. (2023c) Table 175B GPT-3.5 instructions instruction tuning table QA, column-finding,
@@
TableLlama Zhang et al. (2023d) Table 7B LLaMA-2 TableInstruct instruction tuning table QA, RE, entity linking,
@@
TableLLM Zhang et al. (2024d) Table 7B, 13B LLaMA-2 WikiTQ, FeTaQA, instruction tuning table QA, table updating,
@@
LlaSMol Yu et al. (2024) L 6.7B Galactica SMolInstruct instruction tuning QA, classification, regression,
@@
MolCA Liu et al. (2023f) L+G Galactica + PubChem text-graph matching, classification, name conversion,
@@
3D-MoLM Li et al. (2024b) L+G LLaMA-2 + PubChem, 3D-MoIT text-graph matching, QA, regression,
@@
GIT-Mol Liu et al. (2024) L+G+V BERT + PubChem, ChEBI-20 text-graph/image/text classification,
@@
MolGen Fang et al. (2023b) Molecule 406M BART ZINC, NPASS sequence to sequence, molecule generation
@@
PolyNC Qiu et al. (2024) Molecule 220M T5 density functional theory, sequence to sequence classification, regression
@@ 8.9B PMC, MIMIC-III, sentence similarity estimation
ChatDoctor Li et al. (2023f) L 7B LLaMA HealthCareMagic instruction tuning dialogue
@@
AlpaCare Zhang et al. (2023e) L 7B, 13B LLaMA MedInstruct-52k instruction tuning QA, summarization
@@
MUMC Li et al. (2023d) L+V BERT + ROCO, MedICaT, MLM, VQA
@@
CITE Zhang et al. (2023j) L+V BERT + PatchGastric text-image matching, classification
@@
DNABERT-2 Zhou et al. (2023) DNA 110M BERT GRCh38, MLM chromatin profile prediction,
@@
GeoLM Li et al. (2023g) L 110M, 340M BERT OpenStreetMap, MLM, NER, RE, entity typing,
@@
GeoGalactica Lin et al. (2024b) L 30B Galactica geoscience papers, code, next token prediction, QA, knowledge probing,

### v2

@@ Appendix A Summary Tables of Scientific LLMs
CATTS Cachola et al. (2020) L 406M BART SciTLDR sequence to sequence summarization
@@
SciMult Zhang et al. (2023f) L+G 138M MoE MAG, classification, classification, link prediction,
@@
MAmmoTH Yue et al. (2024b) L 7B, 13B, LLaMA-2 MathInstruct instruction tuning QA, MWP
@@
MetaMath Yu et al. (2024c) L 7B, 13B, LLaMA-2 MetaMathQA instruction tuning QA, MWP
@@
MathCoder Wang et al. (2024c) L 7B, 13B, LLaMA-2 MathCodeInstruct instruction tuning QA, MWP
@@
OVM Yu et al. (2024b) L 7B LLaMA-2 GSM8K supervised fine-tuning QA, MWP, quantitative reasoning
@@
MAmmoTH2 Yue et al. (2024c) L 8B LLaMA-3 WebInstruct instruction tuning QA, MWP, quantitative reasoning
@@
     Liang et al. (2023) VQ-VAE TabMWP sequence to sequence geometry problem solving
@@
Table-GPT Li et al. (2024b) Table 175B GPT-3.5 instructions instruction tuning table QA, column-finding,
@@
TableLlama Zhang et al. (2024e) Table 7B LLaMA-2 TableInstruct instruction tuning table QA, RE, entity linking,
@@
TableLLM Zhang et al. (2024f) Table 7B, 13B LLaMA-2 WikiTQ, FeTaQA, instruction tuning table QA, table updating,
@@
PhysBERT Hellert et al. (2024) L 110M BERT arXiv MLM, classification, retrieval,
contrastive learning clustering
@@
Recycle-BERT Kumar et al. (2023) L 110M BERT plastic recycling articles classification QA, classification
@@
LlaSMol Yu et al. (2024a) L 6.7B Galactica SMolInstruct instruction tuning QA, classification, regression,
@@
MolCA Liu et al. (2023e) L+G Galactica + PubChem text-graph matching, classification, name conversion,
@@
3D-MoLM Li et al. (2024c) L+G LLaMA-2 + PubChem, 3D-MoIT text-graph matching, QA, regression,
@@
GIT-Mol Liu et al. (2024a) L+G+V BERT + PubChem, ChEBI-20 text-graph/image/text classification,
@@
Uni-Mol Zhou et al. (2023) Molecule SE(3) ZINC, ChEMBL, 3D position recovery classification, regression,
Transformer RCSB PDB molecule conformation generation,
binding pose prediction
@@
MolGen Fang et al. (2024) Molecule 406M BART ZINC, NPASS sequence to sequence, molecule generation
@@
PolyNC Qiu et al. (2024a) Molecule 220M T5 density functional theory, sequence to sequence classification, regression
@@ 8.9B PMC, MIMIC-III, sentence similarity estimation
ChatDoctor Li et al. (2023e) L 7B LLaMA HealthCareMagic instruction tuning dialogue
@@
AlpaCare Zhang et al. (2023d) L 7B, 13B LLaMA MedInstruct-52k instruction tuning QA, summarization
@@
MMedLM Qiu et al. (2024b) L 7B InternLM MMedC next token prediction QA
1.8B, 7B InternLM2
8B LLaMA-3
@@
Panacea Lin et al. (2024a) L 7B Mistral TrialAlign, next token prediction, summarization, query generation,
TrialInstruct instruction tuning query expansion, trial design,
patient-trial matching
@@
[…]
182M ResNet + PathVQA, Peir Gross, masked image modeling, summarization, image captioning,
[…]
OIA-DDR, CheXpert, VQA, treatment suggestion,
CytoImageNet, ISIC, image captioning mortality prediction
Retinal Fundus,
MIMIC-III, BioNLP,
PubMed
@@
MUMC Li et al. (2023c) L+V BERT + ROCO, MedICaT, MLM, VQA
@@
CITE Zhang et al. (2023i) L+V BERT + PatchGastric text-image matching, classification
@@
Mammo-CLIP Ghosh et al. (2024) L+V BERT + UPMC, VinDr-Mammo text-image matching classification, localization
EfficientNet
@@
DNABERT-2 Zhou et al. (2024a) DNA 110M BERT GRCh38, MLM chromatin profile prediction,
@@
GeoLM Li et al. (2023f) L 110M, 340M BERT OpenStreetMap, MLM, NER, RE, entity typing,
@@
GeoGalactica Lin et al. (2024c) L 30B Galactica geoscience papers, code, next token prediction, QA, knowledge probing,